# Kiss3DGen: Repurposing Image Diffusion Models for 3D Asset Generation

### v1

@@ Abstract
Diffusion models have achieved great success in generating 2D images. However, the quality and generalizability of 3D content generation remain limited. State-of-the-art methods often require large-scale 3D assets for training, which are challenging to collect. In this work, we introduce Kiss3DGen (Keep It Simple and Straightforward in 3D Generation), an efficient framework for generating, editing, and enhancing 3D objects by repurposing a well-trained 2D image diffusion model for 3D generation. Specifically, we fine-tune a diffusion model to generate “3D Bundle Image”, a tiled representation composed of multi-view images and their corresponding normal maps. The normal maps are then used to reconstruct a 3D mesh, and the multi-view images provide texture mapping, resulting in a complete 3D model. This simple method effectively transforms the 3D generation problem into a 2D image generation task, maximizing the utilization of knowledge in pretrained diffusion models. Furthermore, we demonstrate that our Kiss3DGen model is compatible with various diffusion model techniques, enabling advanced features such as 3D editing, mesh and texture enhancement, etc. Through extensive experiments, we demonstrate the effectiveness of our approach, showcasing its ability to produce high-quality 3D models efficiently. Project page: https://ltt-o.github.io/Kiss3dgen.github.io.
@@ 1 Introduction
00footnotetext: This project is supported by Quwan Technology.
@@
Recent research has indicated that 2D diffusion models inherently possess powerful 3D priors that can be utilized for 3D understanding and generation [12, 17]. However, these methods focus on producing 2.5D representations, such as depth maps or normal maps, which provide only a partial view of the 3D structure and are insufficient for full 3D generation. This raises an important question: Can we leverage the expressive priors learned by 2D diffusion models to create complete 3D representations?
@@
Since Kiss3DGen is fundamentally a diffusion model, it is naturally compatible with various diffusion-based techniques. In this work, we demonstrate this by incorporating ControlNet [63] to extend Kiss3DGen’s capabilities, which is termed Kiss3DGen-ControlNet. Given a 3D mesh, Kiss3DGen first renders a 3D Bundle Image, which serves as a condition for ControlNet to perform tasks such as mesh enhancement or editing tasks. Leveraging ControlNet, Kiss3DGen efficiently handles Image-to-3D generation (e.g., enhancing a coarse mesh generated from direct 3D generation), 3D editing (e.g., using specialized ControlNets like Canny), and quality enhancement via ControlNet-Tile, which upgrades low-quality 3D meshes to higher-quality versions.
@@ 2 Related Works
Multi-view generation. Multi-view generation aims to generate multiple viewpoints of an object with a given image or textual prompt. Early efforts in multi-view generation models, such as MVDream [48], have successfully adapted pretrained text-to-image diffusion models [42] to generate object-centric multi-view images. Concurrent studies [27, 47, 50, 34] explore image-conditioned multi-view generation, achieving impressive quality in multi-view outputs. However, they primarily focus on multi-view RGB generation and often overlook the challenges of 3D reconstruction. While some studies [56, 32] explore the separate generation of color and normal maps, another line of research achieves joint generation of both modalities with model context switcher [30, 20]. The generation of normal images profoundly enhances the accuracy and quality of 3D shape formation. However, these approaches significantly modify the network architecture or the input-output patterns of the pretrained models and eliminate textural conditions from their frameworks, thus reducing their effectiveness for general tasks like 3D refinement and editing. To our knowledge, no existing model unifies color and geometry generation with text conditions, a capability we believe is crucial for extending models to various 3D generation tasks, which we have achieved (see Sec. 3).
@@ 3 Proposed Method
Figure 2: The overview of our text-to-3D training and generation framework. In this work, we curate a high-quality text-3D dataset, then train a LoRA [15] layer for text to 3D bundle image (Sec. 3) generation upon a pretrained text-to-image diffusion transformer model with flow matching [25]. After training, our framework generates 3D assets with text condition in two stages: the 3D-Bundle-Image (Sec. 3) generation (Stage I) and the 3D reconstruction (Stage II). In Stage I, we generate 3D bundle image with our Kiss3DGen base model guided by text prompts. In Stage II, we reconstruct the geometry and texture of the 3D asset via LRM [60, 14] or sphere initialization followed by optimization-based mesh refinement and texture projection approach, i.e., ISOMER [56]. Please zoom in for details.
@@ 3.2 Kiss3DGen-ControlNet
To extend the capabilities of Kiss3DGen beyond direct generation, we introduce Kiss3DGen-ControlNet, which incorporates ControlNet [63] to handle a variety of 3D-related tasks such as enhancement, editing, and image-to-3D generation. In this section, we start with 3D enhancement as a basic application and expand into multiple use cases, providing examples of how Kiss3DGen-ControlNet can be applied for the aforementioned tasks. It should be noted that its potential usages are far more extensive than the ones discussed here.
@@
Image-to-3D Generation.  Kiss3DGen-ControlNet also supports Image-to-3D Generation. By using existing methods (e.g., InstantMesh [60]) to generate a coarse mesh from a given 2D image, Kiss3DGen-ControlNet can refine this initial output, transforming it from a low-quality, rough mesh into a high-quality 3D model. In Fig. 3(a, b), we demonstrate an example of image-to-3D generation by reusing our 3D enhancement pipeline, where we further replace one of the RGB views in the rendered 3D Bundle Image with the input image. This two-stage approach allows for the efficient transformation of 2D inputs into detailed 3D objects, utilizing the enhancement capabilities to improve mesh quality significantly.
@@ 4.3 Evaluation
For the evaluation protocol, we assessed alignment with text descriptions and the visual quality of the generated results. For the text-to-3D task, without precise 3D ground truth, we measured alignment using CLIP score [39] by rendering four orthogonal views of each generated object. Additionally, we used Q-Align [55], a large multi-modal model, to evaluate the quality and aesthetics of the rendered images. For the image-to-3D task, we evaluated both 2D visual quality and 3D geometric quality. For the 2D visual evaluation, we rendered novel views from the generation 3D mesh and compared them to the ground truth views using PSNR, SSIM, and LPIPS as metrics. For the 3D geometric evaluation, we focused on comparing the generation 3D meshes to the ground truth. First, we aligned the coordinate systems of the generated and ground truth meshes. Next, we repositioned and rescaled all meshes into a cube of size $[-1,1]^{3}$. We reported Chamfer Distance (CD) and F-Score (FS) at a threshold of 0.1, calculated by uniformly sampling 16K points from the mesh surfaces.
@@
Table 2: Quantitative comparison of text-to-3D generation results after rendering, evaluated in terms of CLIP-score, Quality, and Aesthetic metrics. Our method outperforms 3DTopia [13], Direct2.5 [32], and Hunyuan3D-1.0 [61], achieving higher scores across all metrics, indicating a significant improvement in alignment with textual descriptions and visual quality.
@@ 4.4 Comparison with State-of-the-Art Methods
In this section, we compare our approach with several state-of-the-art methods across three tasks: text-to-multi-view synthesis, text-to-3D generation, and image-to-3D generation. For text-to-multi-view synthesis, we compare with MVDream [48], which introduces a multi-view attention mechanism within its model to facilitate multi-view information interaction. This approach aims to improve consistency across generated views by allowing information to be shared among them. For text-to-3D generation, we compare our method with three recent approaches—3DTopia [13], Direct2.5 [32], and Hunyuan3D-1.0 [61]—that claim capabilities in text-to-3D generation. 3DTopia trains a latent diffusion model to generate Tri-plane representations for 3D object synthesis. Direct2.5 takes a different approach by utilizing two separate diffusion models to generate Normal maps and corresponding RGB maps, which are then used to reconstruct the geometry and apply textures. In contrast, Hunyuan3D-1.0 follows large reconstruction model (LRM) that first generates a single image from text and then expands this into six multi-view images for sparse-view reconstruction. For image-to-3D generation, in addition to the LRM-based Hunyuan3D-1.0, we also compare with CraftsMan [21], a diffusion model-based large-scale image-to-3D generation model, and Unique3D [56], a two stage generation method similar to ours.
Text-to-Multi-View Synthesis. As shown in Tab. 1, we conducted a quantitative evaluation comparing our method to MVDream. Our method demonstrates a substantial improvement over MVDream across all metrics, indicating enhanced consistency and quality across multiple views. Notably, our approach even outperforms the “Real Data” in both “Quality” and “Aesthetic” scores, which may seem counter-intuitive. However, this makes sense considering that our model inherits a lot of knowledge from Flux model which tends to produce high-quality images. Furthermore, even the model trained on a reduced dataset of just 50K samples achieves competitive results, indicating that our approach is highly data-efficient and can yield strong performance even with limited training data. Additionally, we performed qualitative comparisons (see Fig. 4) to further illustrate the superior multi-view coherence and realism achieved by our approach.
Figure 4: Qualitative comparisons with MVDream [48] for text-to-multiview generation. Within the context of text-conditioned multi-view generation, our method produces significantly better results in both text-image alignment and geometric coherence.
@@
Figure 6: Text-to-3D generation comparison between our Base and Doll models (Sec. 4.1). Each model generates different results with different seeds. All images are rendered from 3D mesh.
Figure 7: Qualitative comparison on 3D mesh Enhancement and Editing with MVEdit [3]. Our results (line 2) maintain significantly better consistency with the input mesh in enhancement and better align with the given text condition in editing.
@@
Figure 8: Qualitative comparisons with state-of-the-art methods for image-to-3D generation. Our framework achieves the highest quality 3D mesh, delivering more accurate and realistic texture generation from input images compared to other models.
Image to 3D generation As shown in Tab. 3, we present a quantitative comparison of various image-to-3D generation methods across multiple metrics. Our approach demonstrates superior performance in all evaluated metrics, outperforming other methods in Chamfer Distance (CD), F-Score (FS), PSNR, SSIM, and LPIPS scores. Notably, our model achieves strong results even with a reduced training dataset (Ours-50K), indicating its efficiency and robustness in generating high-quality 3D representations from 2D images. In addition to these quantitative results, Fig. 8 provides qualitative comparisons that highlight the improved texture fidelity, structural coherence, and overall visual quality achieved by our approach.
@@ 5 Concluding Remarks
This work introduces Kiss3DGen, a straightforward yet highly effective approach for a variety of 3D generation tasks. By leveraging knowledge from pretrained 2D diffusion models, it seamlessly integrates with existing techniques like ControlNet. Despite its simplicity, Kiss3DGen excels across tasks such as text-to-3D, image-to-3D, 3D enhancement, and 3D editing. It also demonstrates strong performance even with limited training data, benefiting from the preservation of knowledge in Flux. Additionally, the model generalizes well, enabling the generation of objects beyond its original 3D training set.

### v2

@@ Abstract
footnotetext: * Equal contribution
$\dagger$ Corresponding author
Diffusion models have achieved great success in generating 2D images. However, the quality and generalizability of 3D content generation remain limited. State-of-the-art methods often require large-scale 3D assets for training, which are challenging to collect. In this work, we introduce Kiss3DGen (Keep It Simple and Straightforward in 3D Generation), an efficient framework for generating, editing, and enhancing 3D objects by repurposing a well-trained 2D image diffusion model for 3D generation. Specifically, we fine-tune a diffusion model to generate ”3D Bundle Image”, a tiled representation composed of multi-view images and their corresponding normal maps. The normal maps are then used to reconstruct a 3D mesh, and the multi-view images provide texture mapping, resulting in a complete 3D model. This simple method effectively transforms the 3D generation problem into a 2D image generation task, maximizing the utilization of knowledge in pretrained diffusion models. Furthermore, we demonstrate that our Kiss3DGen model is compatible with various diffusion model techniques, enabling advanced features such as 3D editing, mesh and texture enhancement, etc. Through extensive experiments, we demonstrate the effectiveness of our approach, showcasing its ability to produce high-quality 3D models efficiently. Project page: https://ltt-o.github.io/Kiss3dgen.github.io.
@@ 1 Introduction
footnotetext: This project is supported by Quwan Technology.
@@
Recent research has indicated that 2D diffusion models inherently possess powerful 3D priors that can be utilized for 3D understanding and generation [12, 17]. However, these methods focus on producing 2.5D representations like depth or normal maps, which provide a partial view of the 3D structure and insufficient for full 3D generation. This raises an important question: Can we leverage the expressive priors learned by 2D diffusion models to create complete 3D representations?
@@
Since Kiss3DGen is fundamentally a diffusion model, it is naturally compatible with various diffusion-based techniques. In this work, we demonstrate this by incorporating ControlNet [62] to extend Kiss3DGen’s capabilities, which is termed Kiss3DGen-ControlNet. Given a 3D mesh, Kiss3DGen first renders a 3D Bundle Image, which serves as a condition for ControlNet to perform tasks such as mesh enhancement or editing tasks. Leveraging ControlNet, Kiss3DGen efficiently handles Image-to-3D generation (e.g., enhancing a coarse mesh), 3D editing (e.g., via ControlNet-Canny), and quality enhancement via ControlNet-Tile, which rebuilds low-quality 3D meshes to higher-quality.
@@ 2 Related Works
Multi-view generation. Multi-view generation aims to generate multiple viewpoints of an object with a given image or textual prompt. Early efforts in multi-view generation models, such as MVDream [47], have successfully adapted pretrained text-to-image diffusion models [41] to generate object-centric multi-view images. Concurrent studies [26, 46, 49, 33] explore image-conditioned multi-view generation, achieving impressive quality in multi-view outputs. However, they primarily focus on multi-view RGB generation and often overlook the challenges of 3D reconstruction. While some studies [55, 31] explore the separate generation of color and normal maps, another line of research achieves joint generation of both modalities with model context switcher [29, 20]. The generation of normal images profoundly enhances the accuracy and quality of 3D shape formation. However, these approaches significantly modify the network architecture or the input-output patterns of the pretrained models and discard textural conditions, thus reducing their effectiveness for general tasks like 3D refinement and editing. To our knowledge, no existing model unifies color and geometry generation with text conditions, a capability we believe is crucial for extending models to various 3D generation tasks, which we have achieved (see Sec. 3).
@@ 3 Proposed Method
Figure 2: The overview of our text-to-3D training and generation framework. In this work, we curate a high-quality text-3D dataset, then train a LoRA [15] layer for text to 3D bundle image (Sec. 3) generation upon a pretrained text-to-image diffusion transformer model with flow matching. Our framework generates 3D assets with text condition in two stages: the 3D-Bundle-Image generation (Stage I) and the 3D reconstruction (Stage II). In Stage I, we generate 3D bundle image with our Kiss3DGen base model guided by text prompts. In Stage II, we reconstruct the geometry and texture of the 3D asset via LRM [59, 14] or sphere initialization followed by optimization-based mesh refinement and texture projection approach, i.e., ISOMER [55]. Zoom in for details.
@@ 3.2 Kiss3DGen-ControlNet
To extend the capabilities of Kiss3DGen beyond direct generation, we introduce Kiss3DGen-ControlNet, which incorporates ControlNet [62] to handle a variety of 3D-related tasks such as enhancement, editing, and image-to-3D generation. In this section, we start with 3D enhancement as a basic application and expand into multiple use cases, providing examples of how Kiss3DGen-ControlNet can be applied for the aforementioned tasks. It should be noted that its potential usages are far more extensive.
@@
Image-to-3D Generation.  Kiss3DGen-ControlNet supports Image-to-3D Generation. By using existing methods (e.g., InstantMesh [59]) to generate a coarse mesh from a given 2D image, Kiss3DGen-ControlNet can refine this initial output, transforming it from a low-quality, rough mesh into a high-quality 3D model. In Fig. 3(a, b), we demonstrate an example of image-to-3D generation by reusing our 3D enhancement pipeline, where we further replace one of the RGB views in the rendered 3D Bundle Image with the input image. This two-stage approach allows for the efficient transformation of 2D inputs into detailed 3D objects, utilizing the enhancement capabilities to improve mesh quality.
@@ 4.3 Evaluation
For the evaluation protocol, we assessed alignment with text descriptions and the visual quality of the generated results. For the text-to-3D task, without precise 3D ground truth, we measured alignment using CLIP score [38] by rendering four orthogonal views of each generated object. Also, we used Q-Align [54], a large multi-modal model, to evaluate the quality and aesthetics of the rendered images. For the image-to-3D task, we evaluated both 2D visual quality and 3D geometric quality. For the 2D visual evaluation, we rendered novel views from the generation 3D mesh and compared them to the ground truth views using PSNR, SSIM, and LPIPS as metrics. For the 3D geometric evaluation, we focused on comparing the generation 3D meshes to the ground truth. First, we aligned the coordinate systems of the generated and ground truth meshes. Next, we repositioned and rescaled all meshes into a cube of size $[-1,1]^{3}$. We reported Chamfer Distance (CD) and F-Score (FS) at a threshold of 0.1, calculated by uniformly sampling 16K points from the mesh surfaces.
@@
Table 2: Quantitative comparison of text-to-3D generation results after rendering, evaluated in terms of CLIP-score, Quality, and Aesthetic metrics. Our method outperforms 3DTopia [13], Direct2.5 [31], and Hunyuan3D-1.0 [60] across all metrics, indicating a significant improvement in alignment with textual descriptions and visual quality.
@@ 4.4 Comparison with State-of-the-Art Methods
In this section, we compare our approach with several state-of-the-art methods across three tasks: text-to-multi-view synthesis, text-to-3D generation, and image-to-3D generation. For text-to-multi-view synthesis, we compare with MVDream [47], which introduces a multi-view attention mechanism within its model to facilitate multi-view information interaction. This approach aims to improve consistency across generated views by allowing information to be shared among them. For text-to-3D generation, we compare our method with three recent approaches—3DTopia [13], Direct2.5 [31], and Hunyuan3D-1.0 [60]—that claim capabilities in text-to-3D generation. 3DTopia trains a latent diffusion model to generate Tri-plane representations for 3D object synthesis. Direct2.5 takes a different approach by utilizing two separate diffusion models to generate Normal maps and corresponding RGB maps, which are then used to reconstruct the geometry and apply textures. In contrast, Hunyuan3D-1.0 follows large reconstruction model (LRM) that first generates a single image from text and then expands this into six multi-view images for sparse-view reconstruction. For image-to-3D generation, we compare the Hunyuan3D-1.0, the diffusion-based model CraftsMan [21], and Unique3D [55], a two stage generation method similar to ours.
Text-to-Multi-View Synthesis. As shown in Tab. 1, we conducted a quantitative evaluation comparing our method to MVDream. Our method demonstrates a substantial improvement over MVDream across all metrics, indicating enhanced consistency and quality across multiple views. Notably, our approach outperforms the “Real Data” in both “Quality” and “Aesthetic” scores, which may seem counter-intuitive. However, this makes sense considering that our model inherits a lot of knowledge from Flux model which tends to produce high-quality images. Furthermore, even the model trained on a reduced dataset of just 50K samples achieves competitive results, indicating that our approach is highly data-efficient and can yield strong performance with limited training data. Also, we performed qualitative comparisons in Fig. 4 which illustrate the superior multi-view coherence and realism achieved by our approach.
Figure 4: Qualitative comparisons with MVDream [47] in text-to-multiview generation. In comparison, our method produces significantly better results in both text-image alignment and geometric coherence.
@@
Figure 6: Text-to-3D comparison between our Base and Doll models (Sec. 4.1). Each model generates different results with different seeds. All images are rendered from 3D mesh.
Figure 7: Qualitative comparison on 3D mesh Enhancement and Editing with MVEdit [3]. Our results (line 2) maintain much better consistency with the input mesh in enhancement and better align with the given text in editing.
@@
Figure 8: Qualitative comparisons with state-of-the-art methods for image-to-3D generation. We achieves the highest quality 3D mesh, delivering more accurate and realistic texture generation from input images compared to other models.
Image to 3D generation In Tab. 3, we present a quantitative comparison of various image-to-3D generation methods across multiple metrics. Our approach demonstrates superior performance, outperforming other methods in Chamfer Distance (CD), F-Score (FS), PSNR, SSIM, and LPIPS scores. Notably, our model achieves strong results even with a reduced training dataset (Ours-50K), indicating its efficiency and robustness in generating high-quality 3D representations from 2D images. In addition to these quantitative results, Fig. 8 provides qualitative comparisons that highlight the improved texture fidelity, structural coherence, and overall visual quality achieved by our approach.
@@ 5 Concluding Remarks
We introduces Kiss3DGen, a straightforward yet highly effective approach for a variety of 3D generation tasks. By leveraging knowledge from pretrained 2D diffusion models, it seamlessly integrates with existing techniques like ControlNet. Despite its simplicity, Kiss3DGen excels across tasks such as text-to-3D, image-to-3D, 3D enhancement, and 3D editing. It also demonstrates strong performance even with limited training data, benefiting from the preservation of knowledge in Flux. Additionally, the model generalizes well, enabling the generation of objects beyond its original 3D training set.